OpenStudy (anonymous):

a spherical fish bowl has a radious of 21cm. how many ml of water is needed to fill half the bowl?

OpenStudy (anonymous):

10.5 this might be it but I am not sure

OpenStudy (anonymous):

sorry its 42

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

its 42

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

was I right

OpenStudy (anonymous):

pls give a correct ans.........

OpenStudy (anonymous):

whats ans.......

OpenStudy (anonymous):

one of those are the answer

OpenStudy (anonymous):

where

OpenStudy (anonymous):

10.5 or 42

geerky42 (geerky42):

So you need to find volume of half of spherical fish bowl. You have formula for volume of sphere: \(V = \dfrac{4}{3}\pi r^3\) So half of sphere would be \(V = \dfrac{\dfrac{4}{3}\pi r^3}{2} = \dfrac{4}{6}\pi r^3=\underline{\underline{\underline{\dfrac{2}{3}\pi r^3}}}\) You have \(r=21~cm\). Plug it in half-sphere formula: \(V = \dfrac{2}{3}\pi (21)^3 = \cdots~? ~cm^3\) \(ml\) is same as \(cm^3\). So what is \(\dfrac{2}{3}\pi(21)^3\) ? You need to figure that out by yourself.